OpenStudy (anonymous):

Solve the equation in quadratic formula! 2x^2-9x-2=0

OpenStudy (anonymous):

i got 9\[9\pm \sqrt{65}/4\]

OpenStudy (anonymous):

Then \[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac } }{2a}\] Where a = 2 b = -9 c = 2

OpenStudy (anonymous):

i know how to do that i just got confused at the end

OpenStudy (anonymous):

\[x = \frac{ 9\pm \sqrt{97}}{4}\]

OpenStudy (anonymous):

c = -2 .. sorry my mistake..!

OpenStudy (anonymous):

yeah it should be \[\sqrt{65}\] not\[\sqrt{97}\]

OpenStudy (anonymous):

its \[\sqrt{97}\] since b^2 - 4ac = (-9^2) - (4x 2 x -2) = 81 - (-16) = 97

OpenStudy (anonymous):

ur right i subtracted instead of add. thank you

OpenStudy (anonymous):

Ur Welcome..!! (:

OpenStudy (anonymous):

So you got the answer?

OpenStudy (anonymous):

yes